How many ways can a 4-person subcommittee be selected from a committee of 9 people?
Answer by Alan3354(69443)

The 1st picked is 1 of 9, then 1 of 8, 7, & 6.
= 9*8*7*6 = 3024
But, ABCD is the same as BDCA, so divide by 4*3*2*1 = 24
--> 3024/24 = 126 different ways